# Feature Denoising for Low-Light Instance Segmentation using Weighted Non-Local Blocks

### v1

@@ 5.2 Two-stage experiments
We also compared our method with the two-stage approaches, where the pre-processing employed popular LLIE methods, including EnlightenGAN [7], Zero-DCE++ [8], AGLLNet [14], and RetinexFormer [6]. All of these methods used the weights released by them (for RetinexFormer, we used the weights trained from the LOL [21] dataset). It can be seen from Table 1 that these approaches underperformed compared to using a pretrained model, primarily because the enhanced images were significantly impacted by noise (see Fig. 5, even with the transformer-based method RetinexFormer which reformulated the equation from the Retinex theory to also model noise and reduce the noise amplification. In the example shown, the two-stage methods unsuccessfully generated masks for the giraffe We found that, of all of the LLIE methods used, AGLLNet performed the best. This may be because there model was trained with the synthetic data generated using the same synthetic low-light pipeline as ours.
5.3 Ablation study
We compared our approach with using different NL operations in order to determine which was the most effective method. As seen in Table 2, they all show similar performances, but the embedded Gaussian form shows some improvement compared to the other two. The embedded Gaussian form likely had a higher performance due to its self-attention properties, as described by Wang et al. [13].
| Method | AP | AP | AP |
@@ 5.3 Ablation study
6 Conclusion
This paper introduces a framework for low-light instance segmentation without the need for pre-processing. We enhance popular instance segmentation models by proposing a weighted non-local block for feature denoising, improving performance for smaller objects that would previously most likely be obscured by noise. Experimental results show that our method improves the average precision by +10.0 over a vanilla pretrained Mask R-CNN. Our approach is applicable to various low-light computer vision tasks, including object detection and semantic segmentation, by integrating weighted NL blocks into the backbones of other architectures. Future work should address performance on even smaller objects, enhance other aspects such as the FPN, and assess our method’s efficacy on real low-light datasets.

### v3

@@ IV-B Two-stage experiments
It can be seen from Table II that these approaches underperformed compared to using a pre-trained model, primarily because the contrast-enhanced images were impacted by amplified noise (see Fig. 6), even with the transformer-based method RetinexFormer which attempted to reduce the noise amplification. In the example shown, the two-stage methods made incorrect predictions for the giraffe as the amplified noise confused the model. We found that using AGLLNet for the pre-processing step gave the best results for the two-stage methods. This is likely due to the fact we employed their proposed synthetic low-light pipeline for our synthetic low-light dataset.
IV-C Qualitative assessment on a real low-light dataset
To verify the effectiveness of our method under real low-light conditions, we ran inference on video frames from the BVI-RLV Video Dataset [32], where the low-light videos have 10% of the brightness to the normal-light counterparts. We compared the subjective performance of our wNLB method with the finetuned method, due to a lack of pixel-wise annotations. As shown in Fig. 7, our wNLB method segments more objects and classifies them more accurately than the finetuned model. This confirms that our method can also be applied to instance segmentation tasks under real low-light conditions.
IV-D Non-local operation ablation study
Method | AP | AP50 | AP75 |
@@ IV-D Non-local operation ablation study
We compared our approach with using different NL operations in order to determine which was the most effective denoising algorithm. As seen in Table III, they all show similar performances, but the embedded Gaussian form shows some improvement compared to the other two. The embedded Gaussian form likely had a higher performance due to its self-attention properties, as described by Wang et al. [14].
V Conclusion
This paper introduces a framework for low-light instance segmentation without the need for pre-processing. We enhance popular instance segmentation models by proposing “plug-and-play” weighted non-local blocks (wNLB) for feature denoising. Experimental results show that our method improves the average precision by upto +10.0 over vanilla models pretrained on normal-light images. In particular, our wNLB improves performances for smaller objects obscured by noise that would previously be undetected. Our approach is applicable to various low-light computer vision tasks, including object detection and semantic segmentation, by integrating wNLBs into the backbones of other architectures. Future work should address performance on even smaller objects, improve other aspects such as the mask head for more precise masks, and evaluate our method’s efficacy against real low-light datasets with precise pixel-wise annotations.